1. Which equation shows a slope of -5 and a Y-intercept of (0,1)
2. What is graph for this equation y = -10ver2x + 2
3.What is the graph for this equation y=-2over3 x +2
4Find the x-and y-intercepts of the following equation -3x+4y=-12

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

1. Which equation shows a slope of -5 and a Y-intercept of (0,1)
----------
There are many.
----
2. What is graph for this equation y = -10ver2x + 2
Don't use "over." Use math terms.
-----
3.What is the graph for this equation y=-2over3 x +2
Don't use "over." Use math terms.
-----
4Find the x-and y-intercepts of the following equation -3x+4y=-12
---
At the x-int, y = 0
-3x + 4*0 = -12
-3x = -12
x = 4
---> (4,0)
===========
At the y-int, x = 0
Do it the same way.

Answer by ikleyn(52781)

.

The RULE, the POLICY and the REQUIREMENT of this forum is

    ONE and ONLY ONE problem/question per post.


Thank you.